Hector can shovel snow off the sidewalk in three hours. Oliver can shovel off the same sidewalk in 2 hours. How long will it take them to shovel the snow off the sidewalk if they work together?
Answer by checkley79(3341)

1/3+1/2=1/X FIND THE LCD.(2*3=6)
(1*2+1*3)/2*3=1/X
(2+3)/6=1/X
5/6=1/X CROSS MULTIPLY.
5X=1*6
X=6/5
X=1.2 HOURS WORKING TOGETHER.